# Quantum Entanglement of Anyonic Charges and Emergent Spacetime Geometry

### v1

@@ I Introduction
Intrinsically topologically ordered phases [wen2017, HaldaneNobel, yangbook] with topological entanglement entropy [Wen2006, Kitaev2006] support anyons [Laughlin1983], which may carry fractional electric charge or remain electrically neutral [Feldman2021]. When an electron tunnels into such a system, it can fractionalize into multiple anyons, each obeying nontrivial quantum statistics [Haldane1991] that are neither fermionic nor bosonic. The stability of fractional anyon charges is inherently linked to suppressed fluctuations in the value of the topological entanglement entropy (TEE) [Le_2024].
@@
In this work, we focus on a different facet of anyons: their mutual entanglement. Although fractional charges are often treated as independent quasiparticles, they are in fact quantum mechanically entangled. The mutual information between such excitations is nonzero, even when they are far apart. It reflects non-local correlations in the ground state and serves as a signature of quantum entanglement.
@@
The holographic principle [Hooft1993, Susskind1995] posits that all physical information in a volume can be encoded on its boundary, suggesting that spatial connectivity emerges from quantum correlations. This idea is realized concretely in Anti-de Sitter (AdS) space, where entanglement gives rise to an emergent bulk geometry, i.e., an information-theoretic geometry [Maldacena1997, Witten1998, Ryu2006]. Van Raamsdonk [van2010] further proposed that connected classical spacetimes require sufficient quantum entanglement; reducing this entanglement leads to disconnected geometries, implying that entanglement acts as the “glue” binding spacetime.
@@
This insight motivated Swingle’s proposal [Swingle2012] that tensor networks [orus2014] like the Multiscale Entanglement Renormalization Ansatz (MERA) [Vidal2007, orus2014] serve as discrete models of AdS geometry. MERA uses unitary disentanglers between coarse-graining steps to remove short-range entanglement while preserving long-range correlations, enabling efficient representations of critical states. The hierarchical layers of the network reflect correlations on different length scales, leading to a hyperbolic geometry. Although entanglement-induced geometry does not always need to resemble AdS, AdS provides a particularly transparent and well-studied example.
@@
In parallel, Maldacena and Susskind introduced the ER = EPR conjecture [Maldacena2013], identifying entanglement (EPR pairs [Einstein1935]) with spacetime connections (Einstein-Rosen bridges [Einstein1935_2]). In AdS/CFT, this is realized as entangled black holes on the boundary that are connected by wormholes in the bulk. Crucially, ER = EPR does not depend on conformal symmetry—it extends to a wide class of quantum systems, including non-conformal ones. Even in such cases, as in MERA, the entanglement structure retains a geometric character, supporting the broader principle that spacetime connectivity arises from quantum entanglement.
@@
Connections between entanglement and emergent geometry have also been extensively explored in condensed matter systems. Near quantum critical points, entanglement scaling reveals intriguing patterns reminiscent of holographic duality [Vidal2003, sachdev2011, sachdev2012]. The variance of TEE across topological phase transitions is reflected in the magnitude of the spatial variation of the mutual information [mut_inf2023]. Broader applications of string-theoretic methods in condensed matter physics are reviewed in Refs. [Hartnoll2009, Nastase2017book].
@@
Disordered zigzag graphene nanoribbons [jeong2019, yang2020, yang2021] (ZGNRs) and two-leg electron ladders [Yang2025] exhibit robust long-range entanglement, characterized by a universal value of the TEE that remains invariant under variations in interaction strength and disorder [yang2021]—a hallmark of intrinsic topological order. Both systems host semions charged with $e/2$ with statistical angle $\pi/2$ and belong to the same universality class. Unlike two-leg electron ladders, the entangled fractional charges in ZGNRs are well separated over many lattice spacings, enabling the definition of an emergent space (information geometry) from the mutual information. Similar entanglement structures are expected in other systems that support anyonic excitations.
@@ II Instantons and the Emergence of Topological Charge in ZGNRs
Disordered ZGNRs support topological fractional edge charges. In the clean and non-interacting limit, they behave as a symmetry-protected topological insulator. However, on-site electron interactions break chiral symmetry, allowing a Goldstone–Wilczek-type mechanism to generate a topological charge of one, rather than one-half. In the presence of disorder, these topological states become coupled, giving rise to instantons that carry a fractional charge $e/2$. Atomically precise ZGNRs have been fabricated in the last ten years [Ruffieux2016, Kolmer2020, houtsma2021atomically].
@@ II.1 Hamiltonian
Disordered ZGNRs are modeled using a Hubbard Hamiltonian that includes the interaction term with on-site repulsion $U$ at each site and the disorder term implemented through random on-site potentials $V_{m}$. In the Hartree-Fock mean-field approximation, the Hamiltonian takes the following form.
@@
Here, $t$ denotes the hopping amplitude, $\widehat{c}^{\dagger}_{m,\sigma}$ the fermion creation operator at site $m$ with spin $\sigma$, and $\widehat{n}_{m,\sigma}$ the corresponding number operator. The site index $m$ refers to the positions and the spin index takes values $\sigma=\uparrow$ (spin-up) or $\downarrow$ (spin-down). The sum of $\langle m,n\rangle$ specifies that the hopping is between the nearest neighbors. The second term represents a short-range disorder potential. The magnitude of the potential at each site $V_{m}$ is drawn from a uniform distribution $\Gamma,\Gamma]$. In the continuum limit, the strength of the disorder is characterized by $n_{\mathrm{imp}}\Gamma^{2}$, where $n_{\mathrm{imp}}$ denotes the density of impurities. In the following, we set $n_{\mathrm{imp}}=10\%$ and our results are not sensitive to the precise value of $n_{\mathrm{imp}}$ [Yang2025]. The Hartree–Fock ground state agrees well with the results obtained from the formulation of the matrix product state of the density matrix renormalization group (DMRG) [yang2022] and from bosonization [Yang2025]. The presence of a gap and localization effects suppresses quantum fluctuations, making the Hartree–Fock approximation particularly effective. In this work, we adopt the Hartree–Fock approach, which enables efficient exploration of large systems.
@@ II.2 Breaking of Chiral Symmetry and Goldstone-Wilczek Approach
It is instructive to first consider the disorder-free zigzag ribbon. In the absence of disorder, the non-interacting ZGNR is a symmetry-protected topological insulator in the sense that the gapless edge states exist owing to a sublattice symmetry called chiral symmetry. The chiral transformation is defined as
@@
where $\widehat{\mathsf{a}}_{i}$ and $\widehat{\mathsf{b}}_{i}$ are the electron annihilation operators on the A and B sublattices, respectively. (The operators $\widehat{c}_{i}$ appearing in the Hamiltonian Eq. (1) can be grouped into two types, $\widehat{\mathsf{a}}_{i}$ and $\widehat{\mathsf{b}}_{i}$.) Under the chiral transformation, the non-interacting Hamiltonian (the first term of Eq. (1)) changes sign. Now the axial symmetry transformation of massless Dirac fermions in odd spatial dimensions is
@@
However, this axial symmetry is broken when a mass term is present. A mass term couples $\psi_{R}$ and $\psi_{L}$, and the axial current is no longer conserved, and the continuity equation for the axial currents in one spatial dimension takes the following form [goldstone1981]:
@@
where $\mathcal{M}(x)$ is the mass term that may depend on the coordinates. Goldstone and Wilczek derived an expression for the quantum number of fractional fermions in the background of the topologically nontrival mass term $\mathcal{M}(x)$ (after subtracting the vacuum contribution) in terms of the topological charge [goldstone1981]. The integration of Eq. (5) (in adiabatic limit) between boundary yields ($\rho(x)=j^{0}(x)$)
@@
Now, if the on-site electron repulsion is included, the chiral sublattice symmetry of Eq. (2) is explicitly broken since the density operator does not change under this transformation. We may expect that the edge states of noninteracting ZGNR may be gone. However, it turns out that the edge states persist, but now they are gapped. These gapped edge states can be understood from the point of view of the relation between the fermion number and the topological configuration of the gap parameters analogous to the Goldstone-Wilczek method [chiral-ZGNR]. The combined mirror and time-reversal transformation play crucial roles in characterizing the topological property of the gap parameters. More concretely, one can construct the analogue of anomaly equation as in Eq. (5) for the Hartree-Fock Hamiltonian by considering the time derivative of (total) spatial current using the Heisenberg equation of motion. A careful study of the equation for the current perpendicular to edges shows that the edge occupations of the ZGNR can be identified with the topological charge linked to the solitonic behavior of the gap parameters [chiral-ZGNR].
It is tempting to try to understand the interacting ZGNR in the scheme of the symmetry-based classification of topological insulators/superconductors as proposed in Refs. [tenfold1, tenfold2]. However, it is not clear to what extent such a classification scheme can be applied, since the scheme is based on the condition of the preservation of gapless boundary states against various perturbations that adhere to relevant symmetries.
@@ II.3 Role of Disorder
Up to this point, we have neglected disorder, which can strongly impact the system beyond the symmetry-protected features. In particular, it breaks several symmetries: chiral, mirror, rotational, and translational. The interplay between disorder and electron-electron interactions is highly nontrivial, giving rise to rich and complex physical behavior. Impurities couple the upper and lower zigzag edge states, generating instantons composed of two fractional charges $e/2$, each localized on one of the edges. In the ground state, numerous such pairs are present, with each $e/2$ charge confined to one edge and its partner residing on the opposite edge (see Fig. 1). Upon doping, additional midgap states appear, each carrying a fractional charge [yang2022]. As a result, disorder drives the system from a symmetry-protected topological insulator to an intrinsically topologically ordered phase characterized by a finite topological entanglement entropy [yang2021]. In addition, bosonization shows that fractional charges exhibit semionic statistics such that their fusion yields an electron with ordinary fermionic statistics [Yang2025].
@@
Figure 1: In the disordered ground state of a zigzag ribbon, fractional charge pairs are localized at opposite edges. Yellow dots schematically represent the site-resolved probability density of a fractional charge. It decays rapidly from the edges. While each $e/2$ fractional charge obeys semionic statistics, the composite object formed by a pair, highlighted by the red curve, obeys Fermi statistics [Yang2025]. The number of composite objects scales linearly with the ribbon length.
@@
An edge fractional charge generated by a spin-up instanton can overlap spatially with that from a spin-down instanton, forming a composite with unit electric charge and total spin zero [yang2020]. Such composites are more energetically favorable than individual fractional charges and may give rise to unusual signatures in combined spin susceptibility and transport measurements [yang2022], similar to those observed in polyacetylene [Chung1984].
@@
Several effects are crucial for understanding the robustness of fractional charges. Electron localization plays a significant role in stabilizing fractional charges. Upon weak doping, midgap localized fractional states with $E\approx 0$ exist [yang2022], but they do not spatially overlap each other which is a hallmark of localization [Altshuler, GV2000]. Furthermore, fractional states are isolated from other states of higher energy by an exponential gap in the DOS. This implies that quantum fluctuations have a negligible effect on fractional charges. As the strength of disorder and/or interchain coupling deviates from the universal region in the parameter space, the variance of TEE is expected to grow, which results in crossover phases of quasi-topological order [Le_2024]. In this case, fractional charges are not well defined.
@@ II.4 Non-Local Correlations
Fractional charges are intrinsically linked to non-local correlations. Although the charges themselves are edge localized, their correlations extend into the bulk along the axis joining the two charges (see Fig. 2), as shown in the variation of the number of spin-up occupations, $\delta n_{m,\uparrow}=n_{m,\uparrow}-n^{0}_{m,\uparrow}$, along horizontal carbon lines [Le_2024]. (The quantities $n_{m,\uparrow}$ and $n^{0}_{m,\uparrow}$ represent the occupation numbers in the presence and absence of disorder, respectively.) Abrupt changes in $\delta n_{m,\uparrow}$ are observed when a pair of fractional charges are present at the zigzag edges. Dotted vertical lines indicate the site indices $x$ where these discontinuities occur. Remarkably, sharp variations at the same horizontal positions persist deep in the bulk, far from the physical edges. This behavior signals the presence of non-local correlations, which are observed similarly in the spin-down occupations $\delta n_{m,\downarrow}$ [Le_2024]. These correlations are the origin of the scale-dependent mutual information discussed below. In this context, the separation of a pair of horizontal carbon lines from the system’s edges may be viewed as a renormalization scale, analogous to the depth coordinate in MERA-like tensor network architectures.
Figure 2: (a) Site index for a ZGNR with length $L$ and width $W$. Each carbon site is labeled by $m=(x,y)$, where $x$ denotes the row index and $y$ the column index. Note that each site labeled by $x$ may correspond to different physical positions across horizontal carbon lines, due to their relative displacement within the graphene lattice. Red (blue) sites are labeled by chirality A (B). The opposite zigzag edge sites that form upper and lower boundaries have different chiralities. (b) Left: Each panel shows $\delta n_{m,\uparrow}$, the change in spin-up occupation at site $m$ between disordered and clean cases, plotted along two opposite horizontal carbon chains. Right: Pairs of opposite chains are shown as horizontal dashed lines. The vertical green line corresponds to a vertical dotted line in the left panels; that is, all sites along the green line share the same horizontal index $x$. Fractional charges appear near the two ends of the green line, where $\delta n_{m,\uparrow}$ changes sharply. The Hamiltonian’s parameters are $(U,\Gamma)=(2t,0.1t)$ and the ZGNR geometry is $(L,W)=(120,8)$.

### v2

@@ I Introduction
Intrinsically topologically ordered phases [57, 17, 11] with topological entanglement entropy (TEE) [39, 29] support anyons [32], which may carry fractional electric charge or remain electrically neutral [13]. When an electron tunnels into such a system, it can fractionalize into multiple anyons, each obeying nontrivial quantum statistics [16] that are neither fermionic nor bosonic. The stability of fractional anyon charges is inherently linked to suppressed fluctuations in the value of the TEE [34].
@@
In this work, we focus on a different facet of anyons: their mutual entanglement. Although fractional charges are often treated as independent particles, they are in fact quantum mechanically entangled. The mutual information between such particles is nonzero, even when they are far apart. It reflects non-local correlations in the ground state and serves as a signature of quantum entanglement.
@@
The holographic principle [20, 51] posits that all physical information in a volume can be encoded on its boundary, suggesting that spatial connectivity emerges from quantum correlations. This idea is realized concretely in Anti-de Sitter (AdS) space, where entanglement gives rise to an emergent bulk geometry, i.e., an information-theoretic geometry [41, 59, 46]. Van Raamsdonk [53] further proposed that connected classical spacetimes require sufficient quantum entanglement; reducing this entanglement leads to disconnected geometries, implying that entanglement acts as the “glue” binding spacetime.
@@
This insight motivated Swingle’s proposal [52] that tensor networks such as the multi-scale entanglement renormalization ansatz (MERA) [55, 43] provide a discrete realization of AdS geometry through entanglement renormalization. MERA uses unitary disentanglers between coarse-graining steps to remove short-range entanglement while preserving long-range correlations, enabling efficient representations of critical states. The hierarchical layers of the network reflect correlations on different length scales, leading to a hyperbolic geometry. Although entanglement-induced geometry does not always need to resemble AdS, AdS provides a particularly transparent and well-studied example.
@@
In parallel, Maldacena and Susskind introduced the ER = EPR conjecture [40], identifying entanglement (EPR pairs [6]) with spacetime connections (Einstein-Rosen bridges [7]). In AdS/CFT, this is realized as entangled black holes on the boundary that are connected by wormholes in the bulk. Crucially, ER = EPR does not depend on conformal symmetry, it extends to a wide class of quantum systems, including non-conformal ones. Even in such cases, as in MERA, the entanglement structure retains a geometric character, supporting the broader principle that spacetime connectivity arises from quantum entanglement.
Interestingly, Einstein’s equations follow from the idea that the vacuum state of quantum fields has maximal entanglement in small regions of spacetime [22]. When matter or geometry is slightly perturbed, spacetime responds in such a way that the total entanglement entropy remains unchanged. This requirement uniquely leads to Einstein’s field equations.
@@
Connections between entanglement and emergent geometry have also been extensively explored in condensed matter systems. Near quantum critical points, entanglement scaling reveals intriguing patterns reminiscent of holographic duality [54, 47, 48]. The variance of TEE across topological phase transitions is reflected in the magnitude of the spatial variation of the mutual information [37]. Broader applications of string-theoretic methods in condensed matter physics are reviewed in Refs. [18, 42].
@@
Disordered zigzag graphene nanoribbons [24, 8, 26] (ZGNRs) and two-leg electron ladders [9] exhibit robust long-range entanglement, characterized by a universal value of the TEE that remains invariant under variations in interaction strength and disorder [26]—a hallmark of intrinsic topological order. Both systems host semions charged with $e/2$ with statistical angle $\pi/2$ and belong to the same universality class [9]. In these phases, the ground state already encodes localized semionic degrees of freedom, in contrast to other intrinsically topological phases where anyons appear only as excitations. Unlike two-leg electron ladders, the entangled fractional charges in ZGNRs are well separated over many lattice spacings, enabling the definition of an emergent space (information geometry) from the mutual information. Similar entanglement structures are expected in other systems that support anyonic excitations.
@@ II Instantons and the Emergence of Topological Charge in ZGNRs
Disordered ZGNRs support topological fractional edge charges. In the clean and non-interacting limit, they behave as a symmetry-protected topological insulator. However, on-site electron interactions break chiral symmetry, allowing a Goldstone–Wilczek-type mechanism to generate a topological charge of one, rather than one-half. In the presence of disorder, these topological states become coupled, giving rise to instantons that carry a fractional charge $e/2$. Atomically precise ZGNRs have been fabricated in the last ten years [44, 31, 21].
@@ II.1 Hamiltonian
Disordered ZGNRs are modeled using a Hubbard Hamiltonian that includes the interaction term with on-site repulsion $U$ at each site and the disorder term implemented through random on-site potentials $V_{m}$. In the Hartree-Fock mean-field approximation, the Hamiltonian takes the following form
@@
Here, $t$ denotes the hopping amplitude, $\widehat{c}^{\dagger}_{m,\sigma}$ the fermion creation operator at site $m$ with spin $\sigma$, and $\widehat{n}_{m,\sigma}$ the corresponding number operator. The site index $m$ refers to the positions and the spin index takes values $\sigma=\uparrow$ (spin-up) or $\downarrow$ (spin-down). The sum of $\langle m,n\rangle$ specifies that the hopping is between the nearest neighbors. The second term represents a short-range disorder potential. The magnitude of the potential at each site $V_{m}$ is drawn from a uniform distribution $\Gamma,\Gamma]$. In the continuum limit, the strength of the disorder is characterized by $n_{\mathrm{imp}}\Gamma^{2}$, where $n_{\mathrm{imp}}$ denotes the density of impurities. In the following, we set $n_{\mathrm{imp}}=10\%$ and our results are not sensitive to the precise value of $n_{\mathrm{imp}}$ [8]. The Hartree–Fock ground state agrees well with the results obtained from the formulation of the matrix product state of the density matrix renormalization group (DMRG) [27] and from bosonization [9]. The presence of a gap and localization effects suppresses quantum fluctuations, making the Hartree–Fock approximation particularly effective. In this work, we adopt the Hartree–Fock approach, which enables efficient exploration of large systems.
@@ II.2 Breaking of Chiral Symmetry and Goldstone-Wilczek Approach
It is instructive to first consider the disorder-free zigzag ribbon. In the absence of disorder, the non-interacting ZGNR is a symmetry-protected topological (SPT) insulator in the sense that the gapless edge states exist owing to a sublattice symmetry called chiral symmetry. The chiral transformation is defined as
@@
where $\widehat{\mathsf{a}}_{i}$ and $\widehat{\mathsf{b}}_{i}$ are the electron annihilation operators on the A and B sublattices, respectively. (The operators $\widehat{c}_{i}$ appearing in the Hamiltonian Eq. (1) can be grouped into two types, $\widehat{\mathsf{a}}_{i}$ and $\widehat{\mathsf{b}}_{i}$.) Under the chiral transformation, the non-interacting Hamiltonian (the first term of Eq. (1)) changes sign.
Symmetries can be analyzed more systematically using a low-energy effective theory. In this regime, the continuum limit of the Hubbard model of ZGNR is described by a Dirac Hamiltonian. The axial symmetry transformation of massless Dirac fermions in odd spatial dimensions is
@@
However, this axial symmetry is broken when a mass term is present. A mass term couples $\psi_{R}$ and $\psi_{L}$, and the axial current is no longer conserved, and the continuity equation for the axial currents in one spatial dimension takes the following form [15]:
@@
where $\mathcal{M}(x)$ is the mass term that may depend on coordinates. Goldstone and Wilczek derived an expression for the quantum number of fractional fermions in the background of the topologically nontrivial mass term $\mathcal{M}(x)$ (after subtracting a vacuum contribution) in terms of the topological charge ($Q_{\rm top}$ below) [15]. The integration of Eq. (5) (in the temporal adiabatic limit) between boundary yields ($\rho(x)=j^{0}(x)$)
@@
Now, if the on-site electron repulsion is included, the chiral sublattice symmetry of Eq. (2) is explicitly broken since the density operator does not change under this transformation. We may expect that the edge states of noninteracting ZGNR may be gone. However, it turns out that the edge states persist, but now they are gapped. These gapped edge states can be understood from the point of view of the relation between the fermion number and the topological configuration of the gap parameters analogous to the Goldstone-Wilczek method discussed above [36]. The combined mirror and time-reversal symmetry is essential for determining the system’s topological properties and for maintaining its SPT classification. More concretely, one can construct the analog of anomaly equation as in Eq. (5) for the Hartree-Fock Hamiltonian by considering the time derivative of (total) spatial current using the Heisenberg equation of motion. A careful study of the equation for the current perpendicular to edges shows that the edge occupations of the ZGNR can be identified with the topological charge linked to the behavior of the gap parameters.
The arguments leading to the topologicla charge $Q_{\text{top}}$ in Eq. (6) apply only to symmetry-protected topological insulators (SPTs), which in our case correspond to disorder-free ZGNRs. In this regime, semions are absent: while edge states exist, they carry integer charge $e$, not fractional charge $e/2$ as in polyacetylene [19]. A detailed discussion of these differences can be found in Ref. [9, 36, 10].
It is tempting to try to understand the interacting ZGNR in the scheme of the symmetry-based classification of topological insulators/superconductors as proposed in Refs. [49, 45]. However, it is not clear to what extent such a classification scheme can be applied since the scheme is based on the condition of the preservation of gapless boundary states against various perturbations that adhere to relevant symmetries.
@@ II.3 Role of Disorder
Up to this point, we have neglected disorder, which can strongly impact the system beyond the symmetry-protected features. Disorder may originate from random defects and/or random variations in the hopping parameters. Such disorder breaks several symmetries, including chiral, mirror, rotational, and translational symmetries, and is not confined to the edges of the system. Disorder induces Anderson localization, an emergent property which requires a macroscopic number of impurities. Nonlocal correlations extending across the entire ribbon [34] (to be discussed below) are present only when disorder is distributed throughout the ribbon. The interplay between disorder and electron-electron interactions is highly nontrivial, giving rise to rich and complex physical behavior.
@@
Impurities couple the upper and lower zigzag edge states, generating instantons composed of two fractional charges $e/2$, each localized on one of the edges. In the ground state, numerous such pairs are present, with each $e/2$ charge confined to one edge and its partner residing on the opposite edge (see Fig. 1). Upon doping, additional midgap states appear, each carrying a fractional charge [27]. As a result, disorder drives the system from a symmetry-protected topological insulator to an intrinsically topologically ordered phase characterized by a finite TEE [26]. In addition, bosonization method shows that fractional charges exhibit semionic statistics such that their fusion yields an electron with ordinary fermionic statistics [9].
@@
Figure 1: In the disordered ground state of a zigzag ribbon, fractional charge pairs are localized at opposite edges. Yellow dots schematically represent the site-resolved probability density of a fractional charge. It decays rapidly from the edges. While each $e/2$ fractional charge obeys semionic statistics, the composite object formed by a pair, highlighted by the red curve, obeys Fermi statistics [9]. The number of composite objects scales linearly with the ribbon length.
@@
An edge fractional charge generated by a spin-up instanton can overlap spatially with that from a spin-down instanton, forming a composite with unit electric charge and total spin zero [8]. Such composites are more energetically favorable than individual fractional charges and may give rise to unusual signatures in combined spin susceptibility and transport measurements [27], similar to those observed in polyacetylene [5].
@@
Several effects are crucial for understanding the robustness of fractional charges. Electron localization plays a significant role in stabilizing fractional charges. Upon weak doping, midgap localized fractionalized states with $E\approx 0$ exist [27], but they do not spatially overlap each other which is a hallmark of localization [1, 14]. Furthermore, fractionalized states are isolated from other states of higher energy by an exponential gap in the density of states. This implies that quantum fluctuations have a negligible effect on fractional charges. As the strength of disorder and/or interchain coupling deviates from the universal region in the parameter space, the variance of TEE is expected to grow, which results in crossover phases of quasi-topological order [34]. In this case, fractional charges are not well defined.
@@ II.4 Non-Local Correlations
Fractional charges are intrinsically linked to non-local correlations. Although the charges themselves are edge localized, their correlations extend into the bulk along the axis joining the two charges (see Fig. 2), as shown in the variation of the number of spin-up occupations, $\delta n_{m,\uparrow}=n_{m,\uparrow}-n^{0}_{m,\uparrow}$, along horizontal carbon lines [34]. The quantities $n_{m,\uparrow}$ and $n^{0}_{m,\uparrow}$ represent the occupation numbers in the presence and absence of disorder, respectively. Abrupt changes in $\delta n_{m,\uparrow}$ are observed when a pair of fractional charges are present at the zigzag edges. Dotted vertical lines in the left panels of Fig. 2 (b) indicate the site indices $x$ where these discontinuities occur. Remarkably, sharp variations at the same horizontal positions persist deep in the bulk, far from the physical edges. This behavior signals the presence of non-local correlations, which are observed similarly in the spin-down occupations $\delta n_{m,\downarrow}$ [34]. Nonlocal correlations extending across the entire ribbon (see Fig. 2 (b)) are present only when disorder is distributed throughout the ribbon. These correlations are the origin of the scale-dependent mutual information discussed below.
Figure 2: (a) Site index for a ZGNR with length $L$ and width $W$. Each carbon site is labeled by $m=(x,y)$, where $x$ denotes the row index and $y$ the column index. Note that each site labeled by $x$ may correspond to different physical positions across horizontal carbon lines, due to their relative displacement within the graphene lattice. Red (blue) sites are labeled by chirality A (B). The opposite zigzag edge sites that form upper and lower boundaries have different chiralities. (b) Left: Each panel shows $\delta n_{m,\uparrow}$, the change in spin-up occupation at site $m$ between disordered and clean cases, plotted along two opposite horizontal carbon chains. Right: Pairs of opposite chains are shown as horizontal dashed lines. The vertical green line corresponds to a vertical dotted line in the left panels; that is, all sites along the green line share the same horizontal index $x$. Fractional charges appear near the two ends of the green line, where $\delta n_{m,\uparrow}$ changes sharply. The Hamiltonian’s parameters are $(U,\Gamma)=(2t,0.1t)$ and the ZGNR geometry is specified by $(L,W)=(120,8)$. The data was obtained from a single disorder realization.